Grande03:

Jacqueline works at an aquarium where she keeps a record of the number of fish. The population of fish increases at a rate of 8.3% per year. The following expression represents the number of fish in the aquarium after t years. 21(1.083)^t The expression that reveals the approximate monthly growth rate of the population of fish can be modeled by the expression a^bt . Rounding to the nearest thousandth, determine the values of a and b that reveal the approximate monthly growth rate of the population of fish. a = 1.083 and b = 12 a = 12 and b = 1.083 a = 12 and b = 1.006 a = 1.006 and b = 12

seraphoftheend:

a=12 and b=1.006 im 90% sure

seraphoftheend:

but idk so cold be rong

AZ:

How many months are there in a year?

Grande03:

thats incorrect sera

Grande03:

@AZ 12

seraphoftheend:

12

seraphoftheend:

60 in 12 months

seraphoftheend:

or something like that

seraphoftheend:

az you good

AZ:

The original expression they gave us has 't' but that's in YEARS we now want something to be in months that means we have to divide the RATE by 12 and we multiply t by 12 this website can probably explain it better https://mathbitsnotebook.com/Algebra2/Exponential/EXTimeRelated.html

AZ:

So if we have an expression like \( \Large a(1 + b)^t\) and that's for every year to make it into a month, it would have to become \( \Large a\left(1 + \dfrac{b}{12}\right)^{12t}\)

AZ:

now the 't' in the bottom one is for months so you can calculate it when t is 2 months or 3 but plugging in t = 2 or 3 but in the first expression, if you plug in t = 2 that means 2 YEARS

Grande03:

thank u i got it now

AZ:

What did you get as your final answer? Remember that 1.083 = 1 + 0.083 so you have to divide the 0.083 part by 12

Grande03:

a = 1.006 and b = 12

AZ:

Bingo!